OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 6. What is the value of y when x = 9?

OpenStudy (anonymous):

3 X 6 = 18 18 / 9 = 2 Y=2

OpenStudy (anonymous):

\[y \alpha \frac{1}{x}\] y=k/x find k, by sustituting x and y then using the constant find the value of y

OpenStudy (apoorvk):

so basically, \[y \alpha \frac 1 {2x}\] or \[y = \frac k {2x} \] k=constant now, when x=3, y=6 put x and y in the equation, you find 'k'=36 so, now you have k=36, x=9. can you find the answer now?

OpenStudy (apoorvk):

@thush its 'twice' x